OpenStudy (anonymous):

Please help me simplify the following equation: √(135bc²d³) times √(5b²d)

OpenStudy (anonymous):

\[135^{\frac{ 1 }{ 2 }}b^{\frac{ 1 }{ 2 }}c^{\frac{ 1 }{ 2 }\times2}d^{\frac{ 1 }{ 2 }\times3}\times5^{\frac{ 1 }{ 2 }}b^{\frac{ 1 }{ 2 }\times2}d^{\frac{ 1 }{ 2 }}\]

OpenStudy (anonymous):

\[135^{\frac{ 1 }{ 2 }}b ^{\frac{ 1 }{ 2 }}c^1d ^{\frac{ 3 }{ 2 }}\times5^{\frac{ 1 }{ 2 }}b^1d ^{\frac{ 1 }{ 2 }}\]

OpenStudy (anonymous):

\[(135\times5)^{\frac{ 1 }{ 2 }}b ^{\frac{ 1 }{ 2 }+1}c ^1d ^{\frac{ 3 }{ 2 }+\frac{ 1 }{ 2 }}\]

OpenStudy (anonymous):

\[675^{\frac{ 1 }{ 2 }}b ^{\frac{ 3 }{ 2 }}c^1d^2\]

OpenStudy (anonymous):

\[\sqrt{}=\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

i convert the square root into exponent